OpenStudy (anonymous):

If a man reduces the selling price of a fan from 400 to 380 his loss increases by 20% .What is the cost price of fan?

OpenStudy (anonymous):

500(:

OpenStudy (anonymous):

Let the cost price be X. As per the information given, the loss increases by 20%, i.e. by a factor of 1.2 1.2 (x - 400) = x - 380 1.2x - 480 = x - 380 1.2 x - x = 480 - 380 0.2 x = 100 x = 100 / 0.2 = 500

OpenStudy (anonymous):

CONSIDER THIS WAY SELLING PRICE=80% COST PRICE 380= .8*X X=475 WHAT IS WRONG IN IT?

OpenStudy (anonymous):

@Anyssah

OpenStudy (anonymous):

check the work i wrote above(:

OpenStudy (anonymous):

i got that but y is the 2nd method wrong?

OpenStudy (anonymous):

What do you mean? Another way you can put it i guess is 500-100=400 X -120=380 380+120 =x X =500

OpenStudy (anonymous):

.20 times 500 = 100 you get .20 by moving the decimal 2 points to the left

OpenStudy (anonymous):

at 20% loss selling price is 80% cost price SELLING PRICE=80% COST PRICE 380= .8*X X=475 what is wrong in it

OpenStudy (anonymous):

@Preetha

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

@mark_o.

OpenStudy (anonymous):

@JayDS

OpenStudy (help!!!!):

500